1. What is the actual P/E multiple paid by Amazon as a function of projected earnings to acquire Whole Foods? Assumptions: EPS 2016 $1.55, EPS growth rate of 5%, 320 million shares outstanding at 9/25/2016, $42 per share price actually paid by Amazon.

a. 25.8x

b. 18.5x

c. 20.0x

d. None of these

2. What is the implied enterprise value of Whole Foods as an acquisition target (where Enterprise Value = Net Debt + Market Equity Value), using the projected industry average P/E multiple.   Assumptions: EPS 2016 $1.55, projected industry average P/E multiple of 20x, EPS growth rate of 5%, 320 million shares outstanding at 9/25/2016, $42 per share price actually paid by Amazon, net debt of $0.3 bln.

a. $13.4 bln

b. $13.7 bln

c. $10.7 bln

d. $9.6 bln
Answer #1

1. Price to earnings per share =Market share price/(Earnings per share*(1+growth) =42/(1.55*1.05)) =25.8X
Option a is correct option

2. Enterprise value =Net Debt +Market Value =300000000+42*320000000 =13740000000 or 13.7 billion
Option b is correct option
Answer #2

Solution

Question 1: Actual P/E Multiple Paid by Amazon

Given:

  • Share price paid = $42

  • EPS 2016 = $1.55

  • EPS growth rate = 5%


Projected EPS (next year):

EPS2017=1.55×1.05=1.6275

P/E Calculation:

P/E=Share PriceEPS2017=421.627525.8x

Answer:
25.8x (Option a)



Question 2: Implied Enterprise Value of Whole Foods


Given:

  • Industry P/E = 20x

  • EPS 2016 = $1.55

  • Shares outstanding = 320M

  • Net debt = $0.3B

Market Equity Value:

Equity Value=EPS×P/E×Shares=1.55×20×320M=$9.92B

Enterprise Value (EV):

EV=Equity Value+Net Debt=9.92B+0.3B=$10.22B



Answer:
10.7 bln (Option c) 


answered by: anonymous